OpenStudy (volleyballlover55):

The line is 2/3 and the inital value is 4, what would be the answer is slope form?

OpenStudy (volleyballlover55):

please asnwer quick!

OpenStudy (kgrendel0324):

okay so the initial value is the y-intercept and 2/3 is the slope so y=mx+b

OpenStudy (volleyballlover55):

ok..

OpenStudy (kgrendel0324):

y=2/3x+4

OpenStudy (volleyballlover55):

ofcourse !

OpenStudy (volleyballlover55):

can you help me with a couple more please?

OpenStudy (kgrendel0324):

Yeahh just @ me so i see it

OpenStudy (volleyballlover55):

ok!

OpenStudy (volleyballlover55):

@kgrendel0324